Find the equation of the ellipse that satisfies given conditions:Vertices (±6, 0), foci (±4, 0).

Grade:

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Vertices (±6, 0), foci (±4, 0)
the center (h,k) can be found as
c(h,k) = (f1 + f2)/2 = (v1 + v2a)/2
(h,k) = ( (4,0) + (-4,0))/2 = (0,0)
a is the distance from the vertex to center
a =6 - 0 = 6
c = distance center to foci = 4
c^2 = a^2 - b^2
16 = 36 - b^2
b^2 = 20
x^2/36 + y^2/ 20 = 1
Thanks and regards
Shaik Aasif
askIItians faculty